The time usage of a certain printer cartridge follows a normal probability distribution with mean 58 hours and standard deviation of 183 minutes. Around what time usage guarantee (in hours) should the manufacturer advertise in order to ensure that only 0.1306% of the cartridges fail to meet the guaranteed time usage?


1
Expert's answer
2022-06-13T16:18:52-0400

183 min=3.05 hrs.

"P(X<x)=P(Z<\frac{x-58}{3.05})=0.1306."

"\frac{x-58}{3.05}=-1.12."

"x=58-3.05*1.12=54.58."

The manufacturer should advertise the time usage of 54.58 hours.